HUBBLE SPACE TELESCOPE DAILY REPORT #965 PERIOD COVERED: 0800 EST Daily Status Report as of 246/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.2 Completed WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make six Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of six observations were received. Two images were reported to be underexposed, one observation was saturated, and the others were okay. 1.3 Completed Three Sets of PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) The Planetary Camera was used to make two internal flat calibration images which will be used for the restoration of Earth flat images. They were altered due to changes in quantum efficiency which occurred after decontamination operations. Both observations was received and displayed in the OSS. They were reported to look okay. 1.4 Completed GHRS/2 3584 (Probing the Voids - II) The High Resolution Spectrograph was used to make observations of the active galactic nuclei of Markarian galaxy MARK-335 which is located behind an extra-galactic void. This is to study the nature, abundance, and distribution of Hydrogen I absorbing clouds. The first set of three spectral observations were reported to look okay. The spectral lines in subsequent observations were reported to be smeared. The execution time of these groups suggests that the problems may lie with the on-board Doppler correction. 1.5 Completed Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled, and no problems were reported. 1.6 Completed FGS 4935 (Parallaxes of Hyades Cluster Members - Cycle 3 High) The Fine Guidance Sensors were used to make astrometric observations of star VA200 which is located in the Hyades cluster to obtain a definitive calibration of its population I zero-age main sequence members. The observation appeared to execute with no problems 1.7 Completed FGS 5058 (Extragalactic Astrometry and Astrophysics FT12-23, Part 2 of 2 - Continuation) Fine Guidance Sensor #3 was used to make astrometric observations of of SAO star 2300-683-IN. These observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. The OSS monitored the observations, and assessed the data quality. No problems were reported. 1.8 Completed FOS/BL 4662 (Exploring Triton's UV Spectrum: The First 1800 - 2500 A0 UV Survey) The Faint Object Spectrograph was used to make three observations for the moon Triton between 1800 and 2500 A0. This region has not been observed before. They are looking for CO, NO, and hydrocarbon absorption lines in its spectra. The OSS monitored the observations, and assessed the data quality. No problems were reported. All three observations were reported to look normal. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of PC 5014 ()WF/PC Cycle 3 Calibration: PC Flats Narrow. 2 3.2 Complete Two Sets of FGS 4892 (Duplicity Among Hyades Stars - Cycle 3 High) 3.3 Complete FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) 3.4 Complete PC 4644 (The Nuclei of Nearby S0 and Spiral Galaxies -I A PC Snapshot Imaging Survey) 3.5 Complete Three Sets of PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) 3.6 Complete FOC/96 4546 (Imaging of UV Bright Stars in the Galactic Bulge) 3.7 Complete Four Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.8 Complete Two Sets of GHRS/2 5039 (GHRS Dark Noise Monitoring - Cycle 3) 3.9 Complete FOS/BL 4662 (Exploring Triton's UV Spectrum: The first 1800 - 2500 Ao UV Survey) 3.10 Complete WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) 3.11 Complete HSP/POL 4888 (Remnant Stars in Supernova Remnants - Augmentation - Cycle 3 High) 3.12 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.13 Complete PC 4664 (Astrophysics Study of the Blue Stragglers in 47 Tuanae: Cycle 3 High) 3.14 Complete FOS/BL 4212 (Internal Wavelength Calibration) 3.15 Complete FOS/BL 4260 (Extreme Stellar Winds and Post - Main - Sequence Evolution ) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time.